OpenStudy (anonymous):

Please help I keep getting the wrong answer :/ In a car lift used in a service station, com- pressed air exerts a force on a small piston of circular cross-section having a radius of 4.83 cm. This pressure is transmitted by a liquid to a second piston of radius 20.3 cm. What force must the compressed air exert in order to lift a car weighing 11700 N? Answer in units of N

OpenStudy (anonymous):

aww thanks for the suggestion @PhysicsGuru but that was incorrect.

OpenStudy (anonymous):

Ok the equation for pressure is P = F/A So the top pressure of the 2nd piston is P = 11700 N / pi(0.203m)^2 The pressure on the 1st piston must be the same so 11700 N / pi(0.203m)^2 = Fair / pi(0.0483m)^2 Solve for Fair

OpenStudy (anonymous):

Thank you! I wasn't dividing by pi